OpenStudy (anonymous):

Use the Pythagorean Theorem to find the missing side of the triangle. Round your solution to the nearest tenth, if necessary. A. 20 B. 25 C. 26 D. 30

OpenStudy (anonymous):

Alright I got a bit confused.. ah any takers to help me out? Step by step?

OpenStudy (anonymous):

\[\large c^2 = a^2 + b^2\]

OpenStudy (anonymous):

a and b are the flat edges

OpenStudy (anonymous):

take the square root of the squared version of them to find c

OpenStudy (anonymous):

It doesnt matter if you make a=24 or a = 7.. You get the same answer regardless.

OpenStudy (anonymous):

you have 2 square them and then add them and then take the square root of them

OpenStudy (anonymous):

What ever you decide to make a, then b is the one you didnt choose.

OpenStudy (anonymous):

We will make a = 24 and b = 7\[\large c^2 = 24^2 + 7^2\]

OpenStudy (anonymous):

see if she can figure it from here

OpenStudy (mertsj):

The legs (short sides) are 7 and 24. You can call either one of them a and the other one is b.

OpenStudy (anonymous):

\[\large c^2 = 576 + 49\] Since C is squared we need to get just C.. So take the square root on both sides.\[\large \sqrt{c^2} = \sqrt{576 + 49}\] The square root cancels out the C squared leaving you with just C\[\large C = \sqrt{576 + 49}\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Alright. Well like I was confused on this on the last problem because I kind of figured it was going to be similar to the other question I posted previously to this one considering they're both the Pythagorean Theorem, and I had said that I wrote the steps down for that problem and I was going over it so I knew how to do that for this problem but I was really confused as to how to get the answer for like: \[a ^{2}+5^{2}=8^{2}\] and then how to get \[a ^{2}+25=64\] Would A be considered as 1?

OpenStudy (anonymous):

no, you just simplify and move the 25 over so that you can get a^2 by itself

OpenStudy (mertsj):

No. You solve for a by subtracting 25 from both sides. The you get a^2 = 39. Then take the square root of both sides to get that x = sqrt39

OpenStudy (anonymous):

yurp

OpenStudy (anonymous):

Okay so, so far I did this to the problem: \[c ^{2}+7^{2}=24^{2}\] \[c ^{2}+49=576\] \[c ^{2}=527\] Is that incorrect or is there any way I could do that? But the big question I have about this is now what do I do that I have that final answer?

OpenStudy (anonymous):

to get the final answer*

OpenStudy (mertsj):

No. The legs are 7 and 24. They are a and b. You are being asked to find the hypotenuse which is c.

OpenStudy (anonymous):

Okay so my equation would like this: \[24^{2}+7^{2}=c ^{2}\]?

OpenStudy (anonymous):

look*

OpenStudy (anonymous):

Okay so since I switched that around, I now have: \[527=c ^{2}\] How do I find the square root of that to get my final answer? That is the one I am really confused on now.

OpenStudy (anonymous):

Anyone there?

OpenStudy (phi):

How did you get c^2 = 527?? isn't it 7^2 + 24^2 = c^2 ? 7*7 = 49 24*24= 576 add them together

OpenStudy (anonymous):

That's what c^2 equals so far and then I have to find the square root of 527 and round it to the nearest tenth...

OpenStudy (mertsj):

yes

OpenStudy (phi):

Check your math.

OpenStudy (anonymous):

Wait Mertsj who are you saying yes to? Me or phi?

OpenStudy (mertsj):

Add 49 and 576 again. You will not get 527

OpenStudy (anonymous):

Okay so c\[c ^{2}=625\] but now how do I find the square root of that?

OpenStudy (phi):

You could use a calculator or you could test the answers

OpenStudy (phi):

or you could factor 625 and look for pairs of numbers

OpenStudy (phi):

or you could have memorized it (many people do!)

OpenStudy (phi):

or type sqrt(625) in the google search window